### v1

# Existence of Friedrich-Wintgen Bound States in the Continuum: Cavity with a Thin Waveguide Opening

Abstract
Bound states in the continuum (BICs) are localized states embedded within a continuum of propagating waves. Perturbations that disrupt BICs typically induce ultra-strong resonances, a phenomenon enabling diverse applications in photonics. This work investigates the existence of BICs in two-dimensional electromagnetic cavities coupled to thin waveguides for H-polarized waves. Our focus is on Friedrich-Wintgen BICs (FW-BICs), which arise from destructive interference between two resonant modes and were identified numerically in rectangular cavities with waveguide openings by Lyapina et al. [J. Fluid Mech., 780 (2015), pp. 370–387]. Here, we rigorously establish the existence of FW-BICs in a broader class of cavity geometries by introducing perturbations to the refractive index under regularity constraints. We show that BICs correspond to intersections of two curves derived implicitly from the governing equations constructed via the mode-matching method. Crucially, we prove that such intersections are guaranteed for sufficiently small waveguide widths, provided that two eigenvalues of the cavity cross and the associated eigenfunctions exhibit non-vanishing coupling to the radiation channel at the cavity-waveguide interface. Furthermore, our approach remains applicable for studying the emergence of FW-BICs under parameter-dependent boundary perturbations to the cavity.
1 Introduction
[…]
In this work, we investigate the existence of BICs in a broader class of 2D cavities coupled to semi-infinite waveguides. We consider cavities featuring a line segment on the boundary from which a semi-infinite waveguide extends. A perturbation to the refractive index, analytically dependent on a parameter $\delta$ under specific regularity conditions, is introduced. By reducing the infinite-dimensional linear system derived from the mode-matching method to two governing equations, we show that BICs are characterized as intersections of two curves in the $(\delta,\mu)$-parameter plane, where $\mu$ denotes the square of the free-space wavenumber. Assuming the waveguide width is sufficiently small, we show these intersections are guaranteed to exist, provided that two eigenvalues of the cavity cross as $\delta$ varies and the associated eigenfunctions exhibit non-vanishing coupling to the radiation channel at the cavity-waveguide interface. The main result is formalized in Theorem 5.13. Our method remains applicable for showing the existence of BICs under parameter-dependent boundary perturbations, if the eigenvalues and eigenfunctions exhibit continuous dependence on these perturbations. This continuity condition is satisfied in specific scenarios (cf. Section VII.6.5 in [10]), including scaling transformations of rectangular cavities as explored in [16]. For additional applications of analogous mode-matching techniques in resonance analysis of subwavelength structures, we refer the readers to [14, 15, 33].
This paper is organized as follows. Section 2 outlines the problem setup and three key conditions required for the analysis. In Section 3, we construct the governing infinite-dimensional linear system. Section 4 reformulates the system into a purely real-valued form and reduces it to finite dimensions. Section 5 rigorously establishes the existence of FW-BICs. Numerical experiments validating the theoretical findings are presented in Section 6. Finally, Section 7 briefly discusses the existence of FW-BICs under boundary perturbations, alongside other potential extensions of our approach.
2 Problem formulation and key requirements
As illustrated in Fig. 1, our problem region $\Omega$ is defined as $\Omega:=\Omega_{\mathrm{in}}\cup\Gamma_{h}\cup\Omega_{\mathrm{out}}$, where $\Omega_{\mathrm{in}}$ represents the cavity region with a line segment on its boundary, $\Omega_{\mathrm{out}}$ denotes the waveguide region of width $h$ extended from the line segment, and $\Gamma_{h}$ is the common boundary between these two regions. Letting $\mathbf{o}$ denote the middle point of $\Gamma_{h}$, we establish a rectangular coordinate system in $\mathbb{R}^{2}$ centered at $\mathbf{o}$. Without loss of generality, we always fix the point $\mathbf{o}$ on $\partial\Omega_{\mathrm{in}}$ as the middle point of $\Gamma_{h}$ when $h$ is allowed to decrease in the subsequent sections. Additionally, $\Omega_{B}$ denotes a rectangular region in $\Omega_{\mathrm{in}}$ that $\Gamma_{h}\subset\partial\Omega_{B}$.
$\Omega_{\mathrm{in}}$$n(\mathbf{x})$$\Omega_{\mathrm{out}}$$h$$\Omega_{\mathrm{in}}$$n(\mathbf{x})$$\Omega_{\mathrm{out}}$$h$$\mathbf{o}$$x_{1}$$x_{2}$${\Gamma_{h}}$${\Omega_{B}}$
Figure 1: A cavity with a thin waveguide opening is represented as $\Omega:=\Omega_{\mathrm{in}}\cup\Gamma_{h}\cup\Omega_{\mathrm{out}}$, where $\Omega_{\mathrm{in}}$ represents the cavity region featuring a line segment on its boundary, $\Omega_{\mathrm{out}}$ denotes the waveguide region of width $h$ extended from the line segment vertically, and $\Gamma_{h}$ is the common boundary between these two regions. Additionally, $n(\mathbf{x})$ denotes the refractive index in $\Omega$, and $\Omega_{B}$ denotes a rectangular region in $\Omega_{\mathrm{in}}$ that $\Gamma_{h}\subset\partial\Omega_{B}$.
We also assume $\Omega_{\mathrm{in}}$ is Lipschitz and use $n(\mathbf{x})\in{L^{\infty}(\Omega)}$ to denote the refractive index in $\Omega$ satisfying
(1) {$&n_{0}\leq{n}(\mathbf{x})\leq{n_{1}},\ &&\text{if}\ \mathbf{x}\in\Omega_{\mathrm{in}};\\ &n(\mathbf{x})=1,\ &&\text{if}\ \mathbf{x}\in\Omega_{\mathrm{out}}\cup\Omega_{B},\\ \end{aligned}\right.$
where $n_{0},n_{1}\in\mathbb{R}^{+}$. Letting $\mathcal{P}$ denote the elliptic operator
(2) $\mathcal{P}=-\nabla\cdot\left(\frac{1}{n^{2}}\nabla\right),$
we consider H-polarized electromagnetic waves in $\Omega$ satisfying the Perfect Electric Conductor (PEC) boundary condition:
(3) $-\mathcal{P}{u}+\mu{u}$ $=0\ \mathrm{in}\ \Omega,$
(4) $\partial_{\bm{\nu}}u$ $=0\ \mathrm{on}\ \partial\Omega,$
where $u$ is the only non-zero component of the magnetic field, $\mu$ denotes the square of free space wavenumber and $\bm{\nu}$ is the outward unit normal on $\partial\Omega$. The corresponding eigenvalue problem in $\Omega_{\mathrm{in}}$ is given by
(5) $-\mathcal{P}{\psi}+\lambda{\psi}$ $=0\ \mathrm{in}\ \Omega_{\mathrm{in}},$
(6) $\partial_{\bm{\nu}}\psi$ $=0\ \mathrm{on}\ \partial\Omega_{\mathrm{in}},$
where $(\lambda,\psi)$ denotes an eigenpair.
Let the refractive index $n(\cdot,\delta)$ depend on another parameter $\delta$. When $\delta=0$, we assume the problem (5)–(6) admits a multiple eigenvalue of multiplicity 2 at $\lambda_{M-2}=\lambda_{M-1}$ with corresponding eigenfunctions $\psi_{M-2}$ and $\psi_{M-1}$ for some integer $M\geq 3$. Here, the subscripts denote the indices of the eigenvalues in ascending order. We investigate the existence of Friedrich-Wintgen BICs satisfying (3)–(4) for $\delta\approx 0$ and $\mu\approx\lambda_{M-1}$ when $h\ll 1$, under the following conditions:
  • A.1.
    The refractive index $n(\mathbf{\cdot},\delta)$ depends analytically on $\delta\in[-\delta_{0},\delta_{0}]$ where $\delta_{0}>0$ and satisfies (1) for each $\delta$.
  • A.2.
    When $\delta=0$, $\psi_{M-2}(\mathbf{o})\neq 0$ and $\psi_{M-1}(\mathbf{o})\neq 0$.
  • A.3.
    At $\delta=0$, the eigenvalue functions $\lambda_{M-2}(\delta)$ and $\lambda_{M-1}(\delta)$ cross, resulting in an eigenvalue of multiplicity 2.
A brief discussion of the necessity of these conditions can be found in Remark 5.15.
3 The mode-matching method
In the $\Omega_{\mathrm{in}}$ region, we expand the waves by the eigenfunctions of (5)–(6), while in the $\Omega_{\mathrm{out}}$ region, waves are expanded by outgoing or exponentially decaying waveguide modes. Matching these expansions on $\Gamma_{h}$ through the continuity of traces and the second Green identity leads to an infinite-dimensional linear system. And a non-trivial solution of the homogeneous problem corresponds to a BIC.
3.1 Eigenfunction expansion in the cavity
For each $\delta$, let
(7) $\lambda_{0}(\delta),\ \lambda_{1}(\delta),\ldots\ \text{and}\ \psi_{0}(\cdot,\delta),\ \psi_{1}(\cdot,\delta),\ldots$
denote the eigenvalues and corresponding normalized eigenfunctions of $\mathcal{P}$ in $\Omega_{\mathrm{in}}$ satisfying (5)–(6).
Lemma 3.1.
Under Condition A.1, we can choose a set of normalized eigenfunctions $\{\psi_{m}(\cdot,\delta)\}$ that the mapping $\delta\to\psi_{m}(\cdot,\delta)$ is continuous from $[-\delta_{0},\delta_{0}]$ into $H^{1}(\Omega_{\mathrm{in}})$ for each $m\in\mathbb{N}$, i.e.,
(8) $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{\mathrm{in}})}\to 0\ \text{as}\ \delta_{1}\to\delta_{2}\ \text{for}\ \delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}].$
Furthermore, the associated eigenvalue satisfies
(9) $\left|\frac{1}{\lambda_{m}(\delta_{1})+1}-\frac{1}{\lambda_{m}(\delta_{2})+1}\right|\leq{e^{C_{0}|\delta_{1}-\delta_{2}|}-1},\ \text{for}\ \delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}],$
where $C_{0}>0$ is a constant independent of $m$.
Proof 3.2.
Let $\Phi_{\delta}(\cdot,\cdot)$ denote the corresponding sesquilinear form of (2) for each $\delta$:
(10) $\Phi_{\delta}(u,v):=(\frac{1}{n^{2}(\cdot,\delta)}\nabla{u},\nabla{v})_{\Omega_{\mathrm{in}}},\ u,v{\in}H^{1}(\Omega_{\mathrm{in}}).$
Finding eigenpairs $(\lambda,\psi)$ of (5)–(6) is equivalent to solve
(11) $\Phi_{\delta}(\psi,v)=\lambda(\psi,v)_{\Omega_{\mathrm{in}}},\ \text{for\ each}\ v{\in}H^{1}(\Omega_{\mathrm{in}}).$
Since $n^{2}(\cdot,\delta)$ varies analytically with respect to $\delta$ and $H^{1}(\Omega_{\mathrm{in}})$ is dense in $L^{2}(\Omega_{\mathrm{in}})$, $\Phi_{\delta}$ is a self-adjoint holomorphic family of forms of type (a) (cf. Sections VII.4.2 and VII.4.8 in [10]). Hence, $\psi_{m}(\cdot,\delta)$ can be chosen as a continuous mapping from $[-\delta_{0},\delta_{0}]$ to $L^{2}(\Omega_{\mathrm{in}})$, and its eigenvalue $\lambda_{m}(\delta)$ satisfies (9) by using the results in Section VII.4.7 of [10]. To prove (8), for each $m\in\mathbb{N}$, direct computation gives:
$\Phi_{\delta_{2}}(\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2}),\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2}))$
$=$ $(\lambda_{m}(\delta_{1})\psi_{m}(\cdot,\delta_{1})-\lambda_{m}(\delta_{2})\psi_{m}(\cdot,\delta_{2}),\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2}))_{\Omega_{\mathrm{in}}}$
(12) $+((\frac{1}{n^{2}(\cdot,\delta_{2})}-\frac{1}{n^{2}(\cdot,\delta_{1})})\nabla\psi_{m}(\cdot,\delta_{1}),\nabla(\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})))_{\Omega_{\mathrm{in}}}.$
where the divergence theorem and the boundary conditions $\partial_{\bm{\nu}}\psi_{m}(\cdot,\delta)=0$ on $\partial\Omega_{\mathrm{in}}$ for every $\delta$ have been applied. The vanishing of (12) as $\delta_{1}\to\delta_{2}$ follows from the $L^{2}$-convergence $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{L^{2}(\Omega_{\mathrm{in}})}\to 0$, the analytic dependence of $n^{2}(\cdot,\delta_{1})$ on $\delta_{1}$, and the uniform boundedness of $\|\psi_{m}(\cdot,\delta_{1})\|_{H^{1}(\Omega_{\mathrm{in}})}$.
In the following analysis, we always assume the eigenvalues $\{\lambda_{m}(\delta)\}$ and eigenfunctions $\{\psi_{m}(\cdot,\delta)\}$ depend continuously on $\delta$, and require
(13) $0=\lambda_{0}(0)<\lambda_{1}(0)\leq\cdots.$
The following properties of $\{\lambda_{m}(\delta)\}$ and $\{\psi_{m}(\cdot,\delta)\}$ hold for each $\delta$ (cf. Theorem 4.12 in [18]):
  • (i).
    The eigenfunctions $\{\psi_{m}\}$ can be chosen to be real-valued, and they form a complete orthonormal basis in $L^{2}(\Omega_{\mathrm{in}})$.
  • (ii).
    $\lambda_{0}=0$, $\lambda_{m}>0$ for each $m\in\mathbb{Z}^{+}$ and $\lambda_{m}\to\infty$ as $m\to\infty$.
  • (iii).
    For any $u\in{H}^{1}(\Omega_{\mathrm{in}})$,
    (14) $u=\sum_{m=0}^{\infty}(u,\psi_{m})_{\Omega_{\mathrm{in}}}\psi_{m}\ \text{in}\ H^{1}(\Omega_{\mathrm{in}}),\ \text{and}$
    (15) $\|u\|^{2}_{H^{1}(\Omega_{\mathrm{in}})}\sim\sum_{m=0}^{\infty}(\lambda_{m}+1)|(\psi_{m},u)_{\Omega_{\mathrm{in}}}|^{2}.$
Hence, we assume solutions to (3)–(4) attain the following expansions in $H^{1}(\Omega_{\mathrm{in}})$:
(16) $u=\sum_{m=0}^{\infty}d_{m}\psi_{m},\ \text{where}\ d_{m}=(u,\psi_{m})_{\Omega_{\mathrm{in}}},$
and adopt the $H^{1}(\Omega_{\mathrm{in}})$ norm defined in (15), which is also given by
(17) $\|u\|_{H^{1}(\Omega_{\mathrm{in}})}=\sqrt{\|u\|^{2}_{L^{2}(\Omega_{\mathrm{in}})}+\|\frac{1}{n}\nabla{u}\|^{2}_{L^{2}(\Omega_{\mathrm{in}})}},\ \text{for}\ u{\in}H^{1}(\Omega_{\mathrm{in}}).$
Moreover, (15) implies that the sequence $\{d_{m}\sqrt{\lambda_{m}+1}\}$ belongs to $\ell^{2}$. Due to the conjugate symmetry of equations (3)–(4), we can assume every BIC is real-valued, implying the sequence $\{d_{m}\}$ is also real-valued.
3.2 Mode-expansion in the waveguide
In the waveguide region $\Omega_{\mathrm{out}}:=(0,\infty)\times(-h/2,h/2)$, we apply the method of separation of variables to the following equations:
(18) $\Delta{u}+\mu{u}=$ $0\ \text{in}\ \Omega_{\mathrm{out}},$
(19) $\partial_{\bm{\nu}}{u}=$ $0\ \text{on}\ \partial\Omega_{\mathrm{out}}\backslash\Gamma_{h},$
and $u$ is supposed to be outgoing. This gives the expansions:
(20) $u=\sum_{j=0}^{\infty}b_{j}e^{{\bf i}\alpha_{j}x_{1}}\phi_{j}(x_{2}),$
where $b_{j}=(u,\phi_{j})_{\Gamma_{h}}$,
(21) $\phi_{j}(x_{2})=\left\{\begin{aligned} &\sqrt{\frac{1}{h}},\ &&\text{if}\ j=0;\\ &\sqrt{\frac{2}{h}}\cos\left(\frac{j\pi}{h}x_{2}+\frac{j\pi}{2}\right),\ &&\text{if}\ j\neq 0,\end{aligned}\right.$
$\alpha_{j}=\sqrt{\mu-\left(\frac{j\pi}{h}\right)^{2}}$. We also introduce a set of functions $\{\eta_{j}\}$
(22) $\eta_{j}:=\sqrt{\frac{2}{h}}\sin\left(\frac{j\pi}{h}x_{2}+\frac{j\pi}{2}\right),\ j\in\mathbb{Z}_{+}.$
Both $\{\phi_{j}\}$ and $\{\eta_{j}\}$ form complete orthonormal bases in $L^{2}(\Gamma_{h})$. Moreover, the following expansions hold:
(23) $u|_{\Gamma_{h}}=$ $\sum_{j=0}^{\infty}b_{j}\phi_{j}$ $\text{in}\ H^{1/2}(\Gamma_{h}),$
(24) $\partial_{\bm{\nu}}u|_{\Gamma_{h}}=$ $\sum_{j=0}^{\infty}b_{j}{\bf i}\alpha_{j}\phi_{j}$ $\text{in}\ \widetilde{H}^{-1/2}(\Gamma_{h}).$
3.3 Mode-matching
We apply the second Green identity (cf. Theorem 4.4 in [18]) to $u$ and $\psi_{m}$ in $\Omega_{\mathrm{in}}$ which gives
(25) $(\mathcal{P}u,\psi_{m})_{\Omega_{\mathrm{in}}}-(u,\mathcal{P}\psi_{m})_{\Omega_{\mathrm{in}}}=(u,\partial_{\bm{\nu}}\psi_{m})_{\partial\Omega_{\mathrm{in}}}-(\partial_{\bm{\nu}}u,\psi_{m})_{\partial\Omega_{\mathrm{in}}}.$
Since $u$ and $\psi_{m}$ satisfy (3)–(4) and (5)–(6) respectively, substituting $\partial_{\bm{v}}u$ in (25) with (24) leads to
(26) $(\lambda_{m}-\mu)d_{m}=\sum_{j=0}^{\infty}b_{j}{\bf i}\alpha_{j}(\phi_{j},\psi_{m})_{\Gamma_{h}},\ \text{for\ each}\ m\in\mathbb{N}.$
Meanwhile, matching (16) to the expansions (23) on $\Gamma_{h}$ and taking the $L^{2}(\Gamma_{h})$ inner products with $\phi_{j}$, we obtain
(27) $\sum_{m=0}^{\infty}d_{m}(\psi_{m},\phi_{j})_{\Gamma_{h}}=b_{j},\ \text{for\ each}\ j\in\mathbb{N}.$
Substituting (27) for $b_{j}$ in (26), we arrive at the following infinite-dimensional linear system:
(28) $\left(\begin{bmatrix}\lambda_{0}-\mu&&\\ &\lambda_{1}-\mu&\\ &&\ddots\end{bmatrix}-\sum_{j=0}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{j}\mathbf{v}^{*}_{j}\right)\begin{bmatrix}d_{0}\\ d_{1}\\ \vdots\end{bmatrix}=\mathbf{0},$
where
(29) $\mathbf{v}_{j}:=\begin{bmatrix}(\phi_{j},\psi_{0})_{\Gamma_{h}}&(\phi_{j},\psi_{1})_{\Gamma_{h}}&\cdots\end{bmatrix}^{T}.$
Hence, BICs are characterized as non-trivial solutions $\{d_{m}\}$ to the homogeneous problem (28).
4 Reduction of the system
Given an integer $M\geq 3$, we make the following assumptions when $\delta=0$:
(30) $\lambda_{M-3}(0)<\lambda_{M-2}(0)=\lambda_{M-1}(0)<\lambda_{M}(0).$
Then the following properties regarding the stability of $\{\lambda_{m}(\delta)\}$ hold.
Lemma 4.1.
Under Condition A.1, for every $\epsilon>0$, there exists a closed interval for $\delta$ such that
(31) $\lambda_{m}(\delta)$ $\leq\lambda_{M-3}(0)+\epsilon,\ \forall{m\leq{M-3}},$
(32) $\lambda_{m}(\delta)$ $\geq\lambda_{M}(0)-\epsilon,\ \ \ \ \forall{m\geq{M}}.$
Proof 4.2.
Since only a finite number of eigenvalues exist for $m\leq{M-3}$, (31) follows directly from the continuity of eigenvalues with respect to $\delta$. We focus on (32). For each $m\geq{M}$,
(33) $\lambda_{m}(0)\geq\lambda_{M}(0)>\lambda_{M}(0)-\epsilon.$
If $\lambda_{M}(0)\leq\epsilon$, then (32) holds trivially. Assuming $\lambda_{M}(0)\geq\epsilon$, we obtain
(34) $\frac{1}{\lambda_{m}(0)+1}\leq\frac{1}{\lambda_{M}(0)+1}<\frac{1}{\lambda_{M}(0)+1-\epsilon}.$
By selecting a range for $\delta$ such that
(35) $e^{C_{0}\delta}-1\leq\frac{1}{\lambda_{M}(0)+1-\epsilon}-\frac{1}{\lambda_{M}(0)+1},$
the uniform equicontinuity condition (9) yields the desired result.
Let $\epsilon>0$ be such that
(36) $\epsilon<\frac{1}{3}\min\Big{\{}\lambda_{M-2}(0)-\lambda_{M-3}(0),\lambda_{M}(0)-\lambda_{M-1}(0)\Big{\}}.$
In the following, we assume $\delta$ varies within a range where (31)–(32) hold and
(37) $\lambda_{M-2}(\delta),\lambda_{M-1}(\delta)\in[\lambda_{M-3}(0)+3\epsilon,\lambda_{M}(0)-3\epsilon].$
Furthermore, we restrict $\mu$ to the interval
(38) $\mu\in[\lambda_{M-3}(0)+2\epsilon,\lambda_{M}(0)-2\epsilon].$
Our goal is to reformulate the system (28) into a finite-dimensional real-valued system for $h\ll 1$.
4.1 Decomposition into real and imaginary components
Letting $h\ll 1$ such that $\mu<{\pi^{2}}/{h^{2}}$, only one propagating mode is permitted in the waveguide. Hence we have
(39) $\alpha_{0}>0,\ \text{and}\ {\bf i}\alpha_{j}<0\ \text{for\ each}\ j\in\mathbb{Z}^{+}.$
Since $\{d_{m}\}$ is real-valued, (28) can be decomposed by considering its real and imaginary parts separately:
(40) $\left(\begin{bmatrix}\lambda_{0}-\mu&&\\ &\lambda_{1}-\mu&\\ &&\ddots\end{bmatrix}-\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{j}\mathbf{v}^{*}_{j}\right)\begin{bmatrix}d_{0}\\ d_{1}\\ \vdots\end{bmatrix}=\mathbf{0},$
(41) $(\phi_{0},\sum_{m=0}^{\infty}d_{m}\psi_{m})_{\Gamma_{h}}=0.$
Consequently, both (40) and (41) are real-valued.
4.2 First reduction
By showing a principal submatrix of (40) is invertible, we reduce (40) to an $M$-dimensional linear system in this subsection.
Lemma 4.3.
Assume $\mu<{\pi^{2}}/{h^{2}}$ and the region $R_{h}:=(-h/\pi,0)\times(-h/2,h/2)$ is included in $\Omega_{\mathrm{in}}$. Then the following estimate holds for $u_{1},u_{2}\in{H}^{1}(\Omega_{\mathrm{in}})$:
(42) $\left|\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\phi_{j},u_{2})_{\Gamma_{h}}(u_{1},\phi_{j})_{\Gamma_{h}}\right|{\leq}n_{1}^{2}\left\|\frac{1}{n}{\nabla}u_{1}\right\|_{[L^{2}(R_{h})]^{2}}\left\|\frac{1}{n}{\nabla}u_{2}\right\|_{[L^{2}(R_{h})]^{2}}.$
Proof 4.4.
We introduce the following auxiliary functions $\chi_{j}$ for $j\in\mathbb{Z}^{+}$:
(43) $\chi_{j}(x_{1})=\left\{\begin{aligned} &0,\ &&\text{if}\ x_{1}\leq-\frac{h}{j\pi};\\ &\frac{j\pi}{h}x_{1}+1,\ &&\text{if}\ x_{1}>-\frac{h}{j\pi}.\end{aligned}\right.$
Using the divergence theorem over $R_{h}$ gives
[…]
$=$ $\int_{R_{h}}\partial_{x_{1}}u_{1}(x_{1},x_{2})\phi_{j}(x_{2})\chi_{j}(x_{1})dx_{2}dx_{1}$
(44) $+\int_{R_{h}}\partial_{x_{2}}u_{1}(x_{1},x_{2})\left(-\frac{h}{j\pi}\eta_{j}(x_{2})\chi^{\prime}_{j}(x_{1})\right)dx_{2}dx_{1},$
where $\mathrm{curl}_{2}(f):=(\partial_{x_{2}}f,-\partial_{x_{1}}f)$. Since both $\{\phi_{j}\}$ and $\{\eta_{j}\}$ form complete orthonormal bases in $L^{2}(\Gamma_{h})$, it follows that $\{\sqrt{{\bf i}\alpha_{j}}\phi_{j}\chi_{j}\}$ and $\{-\frac{h}{j\pi}\sqrt{{\bf i}\alpha_{j}}\eta_{j}\chi_{j}^{\prime}\}$ are orthogonal sequences in $L^{2}(R_{h})$. Moreover, their norms are uniformly bounded with respect to $j$ that
(45) $\left\|\sqrt{{\bf i}\alpha_{j}}\phi_{j}\chi_{j}\right\|_{L^{2}(R_{h})}^{2}=$ $\frac{1}{3}\sqrt{1-\frac{h^{2}}{\pi^{2}j^{2}}\mu},$
(46) $\left\|-\frac{h}{j\pi}\sqrt{{\bf i}\alpha_{j}}\eta_{j}\chi_{j}^{\prime}\right\|_{L^{2}(R_{h})}^{2}=$ $\sqrt{1-\frac{h^{2}}{\pi^{2}j^{2}}\mu}.$
Hence, (42) comes from the Cauchy-Schwartz inequality and Bessel’s inequality.
In the following analysis, we always assume $R_{h}\subset{\Omega_{\mathrm{in}}}$. To reduce the system (40)–(41) into a finite-dimensional form, we adopt the following notation:
(47) $\mathbf{A}:=$ $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{0,M-1,j}\mathbf{v}_{0,M-1,j}^{*},$
(48) $\mathbf{B}:=$ $\mathbf{D}\left(\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}\mathbf{v}_{M,\infty,j}^{*}\right)\mathbf{D},$
(49) $\mathbf{V}:=$ $\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}\mathbf{v}_{0,M-1,j}^{*},$
where
$\mathbf{v}_{M_{1},M_{2},j}:=\begin{bmatrix}(\phi_{j},\psi_{M_{1}})_{\Gamma_{h}}\\ \vdots\\ (\phi_{j},\psi_{M_{2}})_{\Gamma_{h}}\end{bmatrix},\ \mathbf{D}:=\begin{bmatrix}\frac{1}{\sqrt{\lambda_{M}-\mu}}&&\\ &\frac{1}{\sqrt{\lambda_{M+1}-\mu}}&\\ &&\ddots\end{bmatrix},$
and $\mathbf{v}_{M_{1},\infty,j}$ denotes the infinite-dimensional vector $\mathbf{v}_{M_{1},M_{2},j}$ as $M_{2}\to\infty$.
Theorem 4.5.
Under Condition A.1 and the assumption $h\ll 1$, the operator $\mathbf{B}$ is uniformly bounded in $\ell^{2}$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, we can obtain
(50) $\|(\mathbf{I}-\mathbf{B})^{-1}\|_{\ell^{2}\times\ell^{2}}\leq 1,$
where $\mathbf{I}$ denotes the identity operator. Additionally, for each $m=0,\ldots,M-1$, the vector $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\phi_{j},\psi_{m})_{\Gamma_{h}}\mathbf{v}_{M,\infty,j}^{*}\mathbf{D}$ belongs to $\ell^{2}$ within these parameter ranges, with its norm bounded by
(51) $\left\|\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\phi_{j},\psi_{m})_{\Gamma_{h}}\mathbf{v}_{M,\infty,j}^{*}\mathbf{D}\right\|_{\ell^{2}}\leq{n_{1}^{2}}\sqrt{1+\frac{{\lambda_{M}(0)+1}}{\epsilon}}\left\|\frac{1}{n}\nabla\psi_{m}\right\|_{[L^{2}(R_{h})]^{2}}.$
Proof 4.6.
Fix $h>0$ sufficiently small that $\mu<\pi^{2}/h^{2}$ and $R_{h}\subset\Omega_{\mathrm{in}}$. We restrict $\delta$ to a small interval that (31)–(32) are satisfied. Let $\{d_{1,m}\}_{m=M}^{\infty}$ and $\{d_{2,m}\}_{m=M}^{\infty}$ denote sequences in $\ell^{2}$. Then we can obtain:
$(\mathbf{B}\{d_{1,m}\},\{d_{2,m}\})_{\ell^{2}}$
$=$ $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\phi_{j},\sum_{m=M}^{\infty}\frac{d_{1,m}}{\sqrt{\lambda_{m}-\mu}}\psi_{m})_{\Gamma_{h}}(\sum_{m=M}^{\infty}\frac{d_{2,m}}{\sqrt{\lambda_{m}-\mu}}\psi_{m},\phi_{j})_{\Gamma_{h}}$
$\leq$ $n_{1}^{2}\left\|\sum_{m=M}^{\infty}\frac{d_{1,m}\psi_{m}}{\sqrt{\lambda_{m}-\mu}}\right\|_{H^{1}(R_{h})}\left\|\sum_{m=M}^{\infty}\frac{d_{2,m}\psi_{m}}{\sqrt{\lambda_{m}-\mu}}\right\|_{H^{1}(R_{h})}$
$\leq$ $n_{1}^{2}\sqrt{\sum_{m=M}^{\infty}\left(1+\frac{\mu+1}{\lambda_{m}-\mu}\right)|d_{1,m}|^{2}}\sqrt{\sum_{m=M}^{\infty}\left(1+\frac{\mu+1}{\lambda_{m}-\mu}\right)|d_{2,m}|^{2}}$
(52) $\leq$ $n_{1}^{2}\left(1+\frac{\lambda_{M}(0)+1}{\epsilon}\right)\|\{d_{1,m}\}\|_{\ell^{2}}\|\{d_{2,m}\}\|_{\ell^{2}},$
where the first inequality follows from Lemma 4.3, and the second inequality follows from our definition of $H^{1}(\Omega_{\mathrm{in}})$ norm in (15). Since $-\mathbf{B}$ is positive, it follows that
(53) $((\mathbf{I}-\mathbf{B})\mathbf{v},\mathbf{v})_{\ell^{2}}\geq\|\mathbf{v}\|_{\ell^{2}}^{2},\ \text{for}\ \mathbf{v}\in\ell^{2}.$
Then the invertibility of $\mathbf{I}-\mathbf{B}$ comes from the lemma of Lax and Milgram (cf. Lemma 2.32 in [18]). Letting $\widetilde{\mathbf{v}}:=(\mathbf{I}-\mathbf{B})\mathbf{v}$, the equation (53) implies
(54) $\|(\mathbf{I}-\mathbf{B})^{-1}\|_{\ell^{2}\times\ell^{2}}\|\widetilde{\mathbf{v}}\|^{2}_{\ell^{2}}\geq(\widetilde{\mathbf{v}},(\mathbf{I}-\mathbf{B})^{-1}\widetilde{\mathbf{v}})_{\ell^{2}}\geq\|(\mathbf{I}-\mathbf{B})^{-1}\widetilde{\mathbf{v}}\|_{\ell^{2}}^{2},$
which leads to (50). Finally, the estimate (51) can be shown following the same procedures presented in (4.6).
Consider the subsystem in (40) satisfied by $[d_{M},d_{M+1},\ldots]^{T}$:
(55) $(\mathbf{I}-\mathbf{B})\mathbf{D}^{-1}\begin{bmatrix}d_{M}\\ d_{M+1}\\ \vdots\end{bmatrix}=\mathbf{V}\begin{bmatrix}d_{0}\\ \vdots\\ d_{M-1}\end{bmatrix}.$
Theorem 4.5 implies $[d_{M},d_{M+1},\ldots]^{T}$ can be represented by $[d_{0},\cdots,d_{M-1}]^{T}$, which leads to a finite dimensions linear system:
(56) $\left(\begin{bmatrix}\lambda_{0}-\mu&&\\ &\ddots&\\ &&\lambda_{M-1}-\mu\end{bmatrix}-\mathbf{A}-\mathbf{V}^{*}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}\right)\begin{bmatrix}d_{0}\\ \vdots\\ d_{M-1}\end{bmatrix}=\mathbf{0};$
(57) $\left(\mathbf{v}_{0,M-1,0}^{*}+\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}\right)\begin{bmatrix}d_{0}\\ \vdots\\ d_{M-1}\end{bmatrix}=0.$
4.3 Further reduction
We further reduce the system (56)–(57) to a $3\times 2$ linear system for the vector $[d_{M-2},d_{M-1}]^{T}$. By evaluating the determinants of two submatrices, we derive two governing equations. The existence of BICs corresponds to determining pairs $(\delta,\mu)$ for which both equations are satisfied. To formalize this, we introduce the following notation:
$\mathbf{C}_{0,0}:=$ $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{0,M-3,j}\mathbf{v}_{0,M-3,j}^{*}$
(58) $+\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{0,M-3,j}\mathbf{v}_{M,\infty,j}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}\mathbf{v}^{*}_{0,M-3,j},$
$\mathbf{C}_{0,1}:=$ $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{0,M-3,j}\mathbf{v}^{*}_{M-2,M-1,j}$
(59) $+\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{0,M-3,j}\mathbf{v}_{M,\infty,j}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}\mathbf{v}^{*}_{M-2,M-1,j},$
(60) $\mathbf{C}_{1,0}:=$ $\mathbf{C}_{0,1}^{*},$
$\mathbf{C}_{1,1}:=$ $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M-2,M-1,j}\mathbf{v}_{M-2,M-1,j}^{*}$
(61) $+\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M-2,M-1,j}\mathbf{v}_{M,\infty,j}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}\mathbf{v}^{*}_{M-2,M-1,j},$
(62) $\mathbf{L}_{0,0}:=$ $\begin{bmatrix}\lambda_{0}-\mu&&\\ &\ddots&\\ &&\lambda_{M-3}-\mu\end{bmatrix},\ \text{and}\ \mathbf{L}_{1,1}:=\begin{bmatrix}\lambda_{M-2}-\mu&\\ &\lambda_{M-1}-\mu\end{bmatrix}.$
Hence the linear system (56) takes the form:
(63) $\left(\begin{bmatrix}\mathbf{L}_{0,0}&\\ &\mathbf{L}_{1,1}\end{bmatrix}-\begin{bmatrix}\mathbf{C}_{0,0}&\mathbf{C}_{0,1}\\ \mathbf{C}_{1,0}&\mathbf{C}_{1,1}\end{bmatrix}\right)\begin{bmatrix}d_{0}\\ \vdots\\ d_{M-1}\end{bmatrix}=\mathbf{0}.$
Lemma 4.7.
Under Condition A.1 and the assumption $h\ll 1$, each entry in the matrices $\mathbf{C}_{0,0},\ \mathbf{C}_{0,1},\ \mathbf{C}_{1,0}$ and $\mathbf{C}_{1,1}$ converges to 00 uniformly as $h\to 0$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38).
Proof 4.8.
We demonstrate this result for each entry in $\mathbf{C}_{0,0}$. Assume $h\ll 1$ such that $\mu<\pi^{2}/h^{2}$ and $R_{h}\subset\Omega_{\mathrm{in}}$. We restrict $\delta$ to a small interval to ensure the validity of (31)–(32). Let $c_{m_{1},m_{2}}$ denote the entry in the $m_{1}$-th row and $m_{2}$-th column of $\mathbf{C}_{0,0}$. Then $c_{m_{1},m_{2}}$ can be expressed as:
$c_{m_{1},m_{2}}=$ $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\phi_{j},\psi_{m_{1}})_{\Gamma_{h}}(\psi_{m_{2}},\phi_{j})_{\Gamma_{h}}$
(64) $+\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\phi_{j},\psi_{m_{1}})_{\Gamma_{h}}\mathbf{v}_{M,\infty,j}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}(\psi_{m_{2}},\phi_{j})_{\Gamma_{h}}.$
Using Lemma 4.3 and Theorem 4.5, we can obtain the bound:
$|c_{m_{1},m_{2}}|\leq$ $n_{1}^{2}\left\|\frac{1}{n}{\nabla}\psi_{m_{1}}\right\|_{[L^{2}(R_{h})]^{2}}\left\|\frac{1}{n}{\nabla}\psi_{m_{2}}\right\|_{[L^{2}(R_{h})]^{2}}$
(65) $+{n_{1}^{4}}\left(1+\frac{{\lambda_{M}(0)+1}}{\epsilon}\right)\left\|\frac{1}{n}\nabla\psi_{m_{1}}\right\|_{[L^{2}(R_{h})]^{2}}\left\|\frac{1}{n}\nabla\psi_{m_{2}}\right\|_{[L^{2}(R_{h})]^{2}}.$
Thus, it’s sufficient to show the uniform convergence:
$\left\|\frac{1}{n}\nabla\psi_{m}(\cdot,\delta)\right\|_{[L^{2}(R_{h})]^{2}}\to 0\ \text{as}\ h\to 0,\ \text{for}\ m=m_{1},m_{2}.$
For fixed $h$, consider the estimate:
$\left|\left\|\frac{1}{n}\nabla\psi_{m}(\cdot,\delta_{1})\right\|_{[L^{2}(R_{h})]^{2}}-\left\|\frac{1}{n}\nabla\psi_{m}(\cdot,\delta_{2})\right\|_{[L^{2}(R_{h})]^{2}}\right|$
$\leq$ $\left\|\frac{1}{n}\left(\nabla\psi_{m}(\cdot,\delta_{1})-\nabla\psi_{m}(\cdot,\delta_{2})\right)\right\|_{[L^{2}(R_{h})]^{2}}$
(66) $\leq$ $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{\mathrm{in}})}.$
[…]
Theorem 4.9.
Under Condition A.1 and the assumption $h\ll 1$, the matrix $\mathbf{L}_{0,0}-\mathbf{C}_{0,0}$ is invertible for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, there exists a constant $C_{1}>1/\epsilon$ such that
(67) $\|(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\|_{2}\leq{C_{1}}.$
Proof 4.10.
Restrict $\delta$ to a sufficiently small interval ensuring the validity of (31)–(32). For $\mu$ satisfying (38), the matrix $\mathbf{L}_{0,0}$ is negative definite with eigenvalues less than $-\epsilon$, which guarantees its invertibility. By Lemma 4.7, the operator norm $\|\mathbf{C}_{0,0}\|_{2}\to 0$ as $h\to 0$. Choose $h\ll 1$ such that $\|\mathbf{C}_{0,0}\|_{2}\leq\epsilon-1/C_{1}$. Applying the Neumann series expansion, we obtain:
$\|(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\|_{2}\leq$ $\|\mathbf{L}_{0,0}^{-1}\|_{2}\|(\mathbf{I}-\mathbf{L}_{0,0}^{-1}\mathbf{C}_{0,0})^{-1}\|_{2}$
$\leq$ $\|\mathbf{L}_{0,0}^{-1}\|_{2}\sum_{j=0}^{\infty}\|\mathbf{L}_{0,0}^{-1}\mathbf{C}_{0,0}\|_{2}^{n}\leq\frac{1}{\epsilon-\|\mathbf{C}_{0,0}\|}\leq{C_{1}}.$
Utilizing Theorem 4.9, the coefficients $[d_{0},\ldots,d_{M-3}]^{T}$ can be determined from
(68) $\begin{bmatrix}d_{0}\\ \vdots\\ d_{M-3}\end{bmatrix}=(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{C}_{0,1}\begin{bmatrix}d_{M-2}\\ d_{M-1}\end{bmatrix}.$
Hence the system (56)–(57) can be further reduced to
(69) $\left(\mathbf{L}_{1,1}-\mathbf{C}_{1,1}-\mathbf{C}_{1,0}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{C}_{0,1}\right)\begin{bmatrix}d_{M-2}\\ d_{M-1}\end{bmatrix}=\mathbf{0};$
$\Big{(}\mathbf{v}_{M-2,M-1,0}^{*}+\mathbf{v}_{0,M-3,0}^{*}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{C}_{0,1}$
(70) $\qquad\qquad+\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}\begin{bmatrix}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{C}_{0,1}\\ \mathbf{I}_{2}\end{bmatrix}\Big{)}\begin{bmatrix}d_{M-2}\\ d_{M-1}\end{bmatrix}=0,$
where $\mathbf{I}_{2}$ denotes the identity operator in $\mathbb{R}^{2}$. Consequently, BICs serve as non-trivial solutions to the system (69)–(70), which can be summarized as follows:
(71) $\begin{bmatrix}\lambda_{M-2}-\mu+a_{0,0}&a_{0,1}\\ a_{1,0}&\lambda_{M-1}-\mu+a_{1,1}\\ a_{2,0}&a_{2,1}\end{bmatrix}\begin{bmatrix}d_{M-2}\\ d_{M-1}\end{bmatrix}=\mathbf{0}.$
Here $a_{0,0}$, $a_{0,1}$, $a_{1,0}$ and $a_{1,1}$ denote the entries of $-\mathbf{C}_{1,1}-\mathbf{C}_{1,0}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{C}_{0,1}$, while $a_{2,0}$ and $a_{2,1}$ correspond to the coefficients in (70). Thus, the existence of BICs is equivalent to finding solutions to the following equations:
(72) $\lambda_{M-2}-\mu+a_{0,0}-a_{0,1}\frac{a_{2,0}}{a_{2,1}}$ $=0,$
(73) $\lambda_{M-1}-\mu+a_{1,1}-a_{1,0}\frac{a_{2,1}}{a_{2,0}}$ $=0,$
which can be derived by computing the determinants of two submatrices from (71).
5 Existence of BICs under analytic medium perturbations
In this section, we introduce the notation of functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$, defined as:
(74) $f_{0}(\delta,\mu,h)=a_{0,0}-a_{0,1}\frac{a_{2,0}}{a_{2,1}},$
(75) $f_{1}(\delta,\mu,h)=a_{1,1}-a_{1,0}\frac{a_{2,1}}{a_{2,0}}.$
Building on Conditions A.1 and A.2 from Section 2 and $h\ll 1$, we establish the following properties of $f_{0}$ and $f_{1}$ for $\delta$ and $\mu$ in small intervals:
  • uniform convergence: $f_{0},f_{1}\to 0$ as $h\to 0$ uniformly;
  • continuity in $\delta$: $f_{0}$ and $f_{1}$ are continuous with respect to $\delta$;
  • Lipschitz continuity in $\mu$: both $f_{0}$ and $f_{1}$ are Lipschitz continuous in $\mu$ with a Lipschitz constant less than 1.
Consequently, each of the equations (72)–(73) defines an implicit function $\mu(\delta)$. BICs emerge as intersections of the graphs of these implicit functions, and their existence is guaranteed under Condition A.3.
5.1 Uniform convergence of $f_{0}$ and $f_{1}$ to 00 as $h\to 0$
We show $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are well-defined and approach 00 as $h\to 0$ uniformly over small intervals of $\delta$ and $\mu$.
Lemma 5.1.
Let $\Omega_{b}\subset{\Omega_{\mathrm{in}}}$ be a region with a $C^{1}$ boundary satisfying $\overline{\Omega_{b}}\subset\Omega_{B}\cup\{x_{1}=0\}$ and $\Gamma_{h}\subset\partial{\Omega_{b}}$. Under Condition A.1, the mapping $\delta\to\psi_{m}(\cdot,\delta)|_{\Omega_{b}}$ is continuous from $[-\delta_{0},\delta_{0}]$ into $H^{3}(\Omega_{b})$ for each $m\in\mathbb{N}$, i.e.,
(76) $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{3}(\Omega_{b})}\to 0\ \text{as}\ \delta_{1}\to\delta_{2},\ \text{for}\ \delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}].$
Proof 5.2.
For each $m\in\mathbb{N}$, observe that $\lambda_{m}$ and $\psi_{m}$ satisfy the eigenvalue problem (5)–(6). From this, we obtain:
(77) $-\Delta\psi_{m}(\mathbf{x},\delta)=\lambda_{m}(\delta)\psi_{m}(\mathbf{x},\delta),\ \text{for}\ \mathbf{x}\in\Omega_{B}.$
The mapping $\delta\to-\Delta\psi_{m}(\cdot,\delta)|_{\Omega_{B}}$ is therefore continuous from $[-\delta_{0},\delta_{0}]$ into $H^{1}(\Omega_{B})$ by using Lemma 3.1. Since $\partial_{\bm{\nu}}\psi_{m}(\cdot,\delta)=0$ on $\partial\Omega_{\mathrm{in}}$, the regularity theory ensures that $\psi_{m}(\cdot,\delta)\in{H^{3}(\Omega_{b})}$ (cf. Theorem 4.18 in [18]). Furthermore, for $\delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}]$, the estimate
$\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{3}(\Omega_{b})}{\leq}$ $C\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{B})}$
(78) $+C\|\Delta\psi_{m}(\cdot,\delta_{1})-\Delta\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{B})}.$
holds for a constant $C>0$. Finally, (76) follows by combining (8) with the continuity of $\delta\to-\Delta\psi_{m}(\cdot,\delta)|_{\Omega_{B}}$ in $H^{1}(\Omega_{B})$.
Theorem 5.3.
Under Conditions A.1, A.2, and the assumption $h\ll 1$, the functions $f_{0}$ and $f_{1}$ are well-defined for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, the uniform convergence
(79) $\lim_{h\to 0}f_{0}(\cdot,\cdot,h)=0,\ \text{and}\ \lim_{h\to 0}f_{1}(\cdot,\cdot,h)=0$
hold.
Proof 5.4.
From Lemma 4.7 and Theorem 4.9, the terms $a_{0,0}$, $a_{0,1}$, $a_{1,0}$ and $a_{1,1}$ are all well-defined and tend to 00 uniformly as $h\to 0$. We focus on showing the uniform boundedness of ratios ${a_{2,0}}/{a_{2,1}}$ and ${a_{2,1}}/{a_{2,0}}$, and (79) follows. From (70), $a_{2,0}$ can be written as
$a_{2,0}=$ $(\psi_{M-2},\phi_{0})_{\Gamma_{h}}+\mathbf{v}_{0,M-3,0}^{*}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{q}$
$+\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}\mathbf{v}_{0,M-3,j}^{*}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{q}$
(80) $+\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}(\psi_{M-2},\phi_{j})_{\Gamma_{h}}.$
Here, $\mathbf{q}$ denotes the finite-dimensional vector:
$\mathbf{q}:=$ $\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\phi_{j},\psi_{M-2})_{\Gamma_{h}}\mathbf{v}_{0,M-3,j}$
(81) $+\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{0,M-3,j}\mathbf{v}^{*}_{M,\infty,j}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}(\Psi_{M-2},\phi_{j})_{\Gamma_{h}}.$
And the following estimate holds by using Lemma 4.3 and Theorem 4.5:
$\|\mathbf{q}\|_{2}\leq$ ${n_{1}^{2}\left\|\frac{1}{n}{\nabla}\psi_{M-2}\right\|_{[L^{2}(R_{h})]^{2}}\left(\sum_{m=0}^{M-3}\left\|\frac{1}{n}{\nabla}\psi_{m}\right\|^{2}_{[L^{2}(R_{h})]^{2}}\right)^{1/2}}$
(82) $+{n_{1}^{4}}\left(1+\frac{{\lambda_{M}(0)+1}}{\epsilon}\right)\left\|\frac{1}{n}\nabla\psi_{M-2}\right\|_{[L^{2}(R_{h})]^{2}}\left(\sum_{m=0}^{M-3}\left\|\frac{1}{n}{\nabla}\psi_{m}\right\|^{2}_{[L^{2}(R_{h})]^{2}}\right)^{1/2}.$
Accordingly, we have
$|a_{2,0}-(\psi_{M-2},\phi_{0})_{\Gamma_{h}}|$
$\leq$ $C_{1}\|\mathbf{v}_{0,M-3,0}^{*}\|_{2}\|\mathbf{q}\|_{2}$
$+C_{1}n_{1}^{2}\sqrt{1+\frac{{\lambda_{M}(0)+1}}{\epsilon}}\left(\sum_{m=0}^{M-3}\left\|\frac{1}{n}{\nabla}\psi_{m}\right\|^{2}_{[L^{2}(R_{h})]^{2}}\right)^{1/2}\|\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}\|_{\ell^{2}}\|\mathbf{q}\|_{2}$
(83) $+n_{1}^{2}\sqrt{1+\frac{{\lambda_{M}(0)+1}}{\epsilon}}\left\|\frac{1}{n}\nabla\psi_{M-2}\right\|_{[L^{2}(R_{h})]^{2}}\|\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}\|_{\ell^{2}},$
where Theorem 4.9 has been used. Consider the sequence $\{d_{0,m}\}_{m=M}^{\infty}\in\ell^{2}$ and analyze its inner product involving $\mathbf{v}^{*}_{M,\infty,0}\mathbf{D}$. This leads to:
$\left|(\mathbf{v}^{*}_{M,\infty,0}\mathbf{D},\{d_{0,m}\})_{\ell^{2}}\right|=$ $\left|(\phi_{0},\sum_{m=M}^{\infty}\frac{d_{0,m}}{\sqrt{\lambda_{m}-\mu}}\psi_{m})_{\Gamma_{h}}\right|$
$\leq$ $\left\|\sum_{m=M}^{\infty}\frac{d_{0,m}}{\sqrt{\lambda_{m}-\mu}}\psi_{m}\right\|_{L^{2}(\Gamma_{h})}{\leq}C\left\|\sum_{m=M}^{\infty}\frac{d_{0,m}}{\sqrt{\lambda_{m}-\mu}}\psi_{m}\right\|_{H^{1}(\Omega_{\mathrm{in}})}$
(84) $\leq$ $Cn_{1}^{2}\sqrt{1+\frac{{\lambda_{M}(0)+1}}{\epsilon}}\|\{d_{0,m}\}\|_{\ell^{2}},$
where $C$ is a constant independent of $h$. Consequently, both the norms of $\mathbf{v}_{0,M-3,0}^{*}$ and $\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}$ are uniformly bounded. For $h\ll 1$, the region $R_{h}$ lies within $\Omega_{b}$ as defined in Lemma 5.1. By the general Sobolev inequalities (cf. Section 5 in [3]), (76) gives the continuity of the mapping $\delta\to\psi_{m}(\cdot,\delta)|_{\overline{\Omega_{b}}}$ in $C^{1,\gamma}$ for $10<\gamma<$ 1. Thus the following asymptotic expansions hold:
(85) $(\psi_{M-2}(\cdot,\delta),\phi_{0})_{\Gamma_{h}}=$ $\psi_{M-2}(\mathbf{o},\delta)\sqrt{h}+\mathcal{O}(h^{3/2}),$
(86) $\|\mathbf{q}\|_{2}=$ $\mathcal{O}(h^{2}),$
(87) $\|\nabla\psi_{M-2}(\cdot,\delta)\|_{[L^{2}(R_{h})]}=$ $\mathcal{O}(h),$
(88) $\left(\sum_{m=0}^{M-3}\left\|\frac{1}{n}{\nabla}\psi_{m}\right\|^{2}_{[L^{2}(R_{h})]^{2}}\right)^{1/2}=$ $\mathcal{O}(h),$
where the constants implicit in the $\mathcal{O}-$notation are independent of $\delta$. From these results, the leading behavior of $a_{2,0}$ is given by:
(89) $a_{2,0}=\psi_{M-2}(\mathbf{o},\delta)\sqrt{h}+\mathcal{O}(h),$
and analogously,
(90) $a_{2,1}=\psi_{M-1}(\mathbf{o},\delta)\sqrt{h}+\mathcal{O}(h),$
which ensures the uniform boundedness of the ratios ${a_{2,0}}/{a_{2,1}}$ and ${a_{2,1}}/{a_{2,0}}$ for all $\delta$ in a sufficiently small interval.
5.2 Continuity in $\delta$ and Lipschitz continuity in $\mu$ for $f_{0}$ and $f_{1}$
In the following analysis, we establish the continuity of the functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ in the variable $\delta$, as well as their Lipschitz continuity in $\mu$ with a Lipschitz constant less than $1$.
Lemma 5.5.
Let $\mathbf{T}(\delta,\mu)$ denote an operator in $\ell^{2}$ that
(91) $\mathbf{T}(\delta,\mu)=\begin{bmatrix}\frac{\sqrt{1+\lambda_{M}(\delta)}}{\sqrt{\lambda_{M}(\delta)-\mu}}&&\\ &\frac{\sqrt{1+\lambda_{M+1}(\delta)}}{\sqrt{\lambda_{M+1}(\delta)-\mu}}&\\ &&\ddots\end{bmatrix}.$
Under Condition A.1, $\mathbf{T}(\delta,\mu)$ is continues in $\delta$ and $\mu$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Moreover, there is a constant $C_{2}$ irrelevant to $\delta$ that
(92) $\|\mathbf{T}(\delta,\mu_{1})-\mathbf{T}(\delta,\mu_{2})\|_{\ell^{2}\times\ell^{2}}\leq{C_{2}}|\mu_{1}-\mu_{2}|.$
Proof 5.6.
Given $m\geq{M}$, each element in $\mathbf{T}(\delta_{1},\mu)-\mathbf{T}(\delta_{2},\mu)$ satisfies
$\left|\frac{\sqrt{1+\lambda_{m}(\delta_{1})}}{\sqrt{\lambda_{m}(\delta_{1})-\mu}}-\frac{\sqrt{1+\lambda_{m}(\delta_{2})}}{\sqrt{\lambda_{m}(\delta_{2})-\mu}}\right|$
$=$ $\left|\left(\sqrt{1+\frac{\mu+1}{\lambda_{m}(\delta_{1})-\mu}}+\sqrt{1+\frac{\mu+1}{\lambda_{m}(\delta_{2})-\mu}}\right)^{-1}\left(\frac{\mu+1}{\lambda_{m}(\delta_{1})-\mu}-\frac{\mu+1}{\lambda_{m}(\delta_{2})-\mu}\right)\right|$
$\leq$ $\frac{1}{2}(\lambda_{M}(0)+1)\left(1+\frac{\lambda_{M}(0)+1}{\epsilon}\right)^{2}\left(e^{C_{0}|\delta_{1}-\delta_{2}|}-1\right),$
which tends to 00 uniformly. Then the continuity of $\mathbf{T}(\delta,\mu)$ with respect to $\delta$ follows. Similarly, each element in $\mathbf{T}(\delta,\mu_{1})-\mathbf{T}(\delta,\mu_{2})$ satisfies
$\left|\frac{\sqrt{1+\lambda_{m}(\delta)}}{\sqrt{\lambda_{m}(\delta)-\mu_{1}}}-\frac{\sqrt{1+\lambda_{m}(\delta)}}{\sqrt{\lambda_{m}(\delta)-\mu_{2}}}\right|$
$=$ $\sqrt{1+\frac{\mu_{1}+1}{\lambda_{m}(\delta)-\mu_{1}}}\left|\frac{\mu_{1}-\mu_{2}}{\sqrt{\lambda_{m}(\delta)-\mu_{2}}(\sqrt{\lambda_{m}(\delta)-\mu_{1}}+\sqrt{\lambda_{m}(\delta)-\mu_{2}})}\right|$
$\leq$ $\frac{1}{2\epsilon}\sqrt{1+\frac{\lambda_{M}(0)+1}{\epsilon}}|\mu_{1}-\mu_{2}|,$
which gives (92).
To study the continuity of the operator $(\mathbf{I}-\mathbf{B})^{-1}$ in $\ell^{2}$ with respect to $\delta$, we say an operator $\mathbf{G}(\delta)$ in $\ell^{2}$ is strongly continuous on $[\delta_{a},\delta_{b}]\subset\mathbb{R}$ if for every $\mathbf{d}\in\ell^{2}$
(93) $\left\|\mathbf{G}(\delta_{1})\mathbf{d}-\mathbf{G}(\delta_{2})\mathbf{d}\right\|_{\ell^{2}}\to 0\ \text{as}\ \delta_{1}\to\delta_{2},\ \text{for}\ \delta_{1},\delta_{2}\in[\delta_{a},\delta_{b}],$
as defined in [10, p. 152].
Lemma 5.7.
Let $\mathbf{G}(\delta)$ denote an invertible, self-adjoint operator in $\ell^{2}$ that is strongly continuous with respect to $\delta\in[\delta_{a},\delta_{b}]$, where $\delta_{a},\delta_{b}\in\mathbb{R}$. If its inverse $\mathbf{G}^{-1}(\delta)$ is uniformly bounded on $[\delta_{a},\delta_{b}]$, then $\mathbf{G}^{-1}(\delta)$ is also strongly continuous.
Proof 5.8.
Given any uniform bounded vector $\mathbf{v}(\delta)\in\ell^{2}$ and a constant vector $\mathbf{q}\in\ell^{2}$, (93) implies
(94) $(\mathbf{v}(\delta_{1}),\mathbf{G}(\delta_{1})\mathbf{q}-\mathbf{G}(\delta_{2})\mathbf{q})_{\ell^{2}}\to 0\ \text{as}\ \delta_{1}\to\delta_{2}.$
Define $\widetilde{\mathbf{q}}:=\mathbf{G}(\delta_{2})\mathbf{q}$ and $\mathbf{v}(\delta_{1}):=\mathbf{G}^{-1}(\delta_{1})(\mathbf{G}^{-1}(\delta_{2})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{1})\widetilde{\mathbf{q}})$. Substituting into (94) and using the self-adjointness of $\mathbf{G}(\delta_{1})$ and uniform boundedness of $\mathbf{G}^{-1}(\delta_{1})$, we obtain
(95) $(\mathbf{G}^{-1}(\delta_{2})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{1})\widetilde{\mathbf{q}},\mathbf{G}^{-1}(\delta_{2})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{1})\widetilde{\mathbf{q}})_{\ell^{2}}\to 0\ \text{as}\ \delta_{1}\to\delta_{2},$
which implies $\mathbf{G}^{-1}(\delta)$ is strongly continuous.
Lemma 5.9.
Let $\mathbf{H}(\delta,\mu)$ denote an operator in $\ell^{2}$ defined by
(96) $\mathbf{H}(\delta,\mu):=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(k)\mathbf{w}_{j}(\delta)\mathbf{w}_{j}^{*}(\delta),$
where the vectors $\mathbf{w}_{j}(\delta)$ are given by
(97) $\mathbf{w}_{j}(\delta):=\begin{bmatrix}(\phi_{j},\frac{\psi_{M}(\cdot,\delta)}{\sqrt{1+\lambda_{M}(\delta)}})_{\Gamma_{h}}&(\phi_{j},\frac{\psi_{M+1}(\cdot,\delta)}{\sqrt{1+\lambda_{M+1}(\delta)}})_{\Gamma_{h}}&\cdots\end{bmatrix}^{T}.$
Under Condition A.1 and the assumption $h\ll 1$, the operator $\mathbf{H}(\delta,\mu)$ is strongly continuous in $\delta$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, we have
(98) $\|\mathbf{H}(\delta,\mu_{1})-\mathbf{H}(\delta,\mu_{2})\|_{\ell^{2}\times\ell^{2}}\leq{C_{3}}|\mu_{1}-\mu_{2}|,$
where $C_{3}$ is a constant independent of $\delta$.
Proof 5.10.
We consider the operator $\widetilde{\mathbf{H}}(\delta,\mu)$ defined as
(99) $\widetilde{\mathbf{H}}(\delta,\mu):=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)\widetilde{\mathbf{w}}_{j}(\delta)\widetilde{\mathbf{w}}_{j}^{*}(\delta),$
where
(100) $\widetilde{\mathbf{w}}_{j}(\delta):=\begin{bmatrix}(\phi_{j},\frac{\psi_{0}(\cdot,\delta)}{\sqrt{1+\lambda_{0}(\delta)}})_{\Gamma_{h}}&(\phi_{j},\frac{\psi_{1}(\cdot,\delta)}{\sqrt{1+\lambda_{1}(\delta)}})_{\Gamma_{h}}&\cdots\end{bmatrix}^{T}.$
Accordingly, $\mathbf{H}(\delta,\mu)$ serves as a restriction of $\widetilde{\mathbf{H}}(\delta,\mu)$. Let $\ell^{2}_{c}$ denote a subspace of $\ell^{2}$ defined by
(101) $\ell^{2}_{c}:=\left\{\{d_{m}\}\in\ell^{2}:\exists{m_{0}\in\mathbb{N}}\ s.t.\ d_{m}=0\ \text{if}\ m{\geq}m_{0}\right\},$
which is dense in $\ell^{2}$. Given $\{\widetilde{d}_{0,m}\}\in\ell^{2}_{c}$, we first show the continuity of $\mathbf{d}(\delta):=\widetilde{\mathbf{H}}(\delta,\mu)\{\widetilde{d}_{0,m}\}$ in $\ell^{2}$. For any $\{\widetilde{d}_{1,m}\}\in\ell^{2}_{c}$, the inner product $(\mathbf{d}(\delta_{1}),\{\widetilde{d}_{1,m}\})_{\ell^{2}}$ converges to $(\mathbf{d}(\delta_{2}),\{\widetilde{d}_{1,m}\})_{\ell^{2}}$ as $\delta_{1}\to\delta_{2}$ by Lemma 3.1. Hence, the Banach-Steinhaus Theorem (cf. [20]) implies the weak convergence of $\mathbf{d}(\delta_{1})$ to $\mathbf{d}(\delta_{2})$. Let $\{d_{0,m}\}\in\ell^{2}$ with $\|\{d_{0,m}\}\|_{\ell^{2}}=1$ such that
$\|\mathbf{d}(\delta_{2})\|_{\ell^{2}}$ $=(\mathbf{d}(\delta_{2}),\{d_{0,m}\})_{\ell^{2}}$
(102) $=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}})_{\Gamma_{h}}(\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}},\phi_{j})_{\Gamma_{h}}.$
Since $\{{\psi_{m}(\cdot,\delta)}/{\sqrt{1+\lambda_{m}(\delta)}}\}$ forms a complete orthonormal basis in $H^{1}(\Omega_{\mathrm{in}})$, there exists $\{d_{1,m}\}\in\ell^{2}$ such that $\|\{d_{1,m}\}\|_{\ell^{2}}=1$ and
(103) $\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}}\Bigg{|}_{\Gamma_{h}}=\sum_{m=0}^{\infty}\frac{d_{1,m}\psi_{m}(\cdot,\delta_{1})}{\sqrt{1+\lambda_{m}(\delta_{1})}}\Bigg{|}_{\Gamma_{h}}.$
Hence, we have
$\|\mathbf{d}(\delta_{2})\|_{\ell^{2}}-\|\mathbf{d}(\delta_{1})\|_{\ell^{2}}$ $\leq\|\mathbf{d}(\delta_{2})\|_{\ell^{2}}-(\mathbf{d}({\delta_{1}}),\{d_{1,m}\})_{\ell^{2}}$
(104) ${\leq}n_{1}^{2}$ $\left\|\frac{1}{n(\cdot,\delta_{1})}\nabla\left(\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}}-\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{1})}{\sqrt{1+\lambda_{m}(\delta_{1})}}\right)\right\|_{[L^{2}(R_{h})]^{2}}$
by Lemma 4.3. Furthermore, (104) becomes infinitesimally small as $\delta_{1}\to\delta_{2}$ by Lemma 3.1, since $\{\widetilde{d}_{0,m}\}\in\ell^{2}_{c}$. Exchanging $\delta_{1}$ and $\delta_{2}$ in (104) yields
(105) $\|\mathbf{d}(\delta_{1})\|_{\ell^{2}}\to\|\mathbf{d}(\delta_{2})\|_{\ell^{2}},\ \text{as}\ \delta_{1}\to\delta_{2},$
which implies the strong convergence of $\mathbf{d}(\delta_{1})$ to $\mathbf{d}(\delta_{2})$ (cf. Exercise 2.11 in [18]). Therefore $\widetilde{\mathbf{H}}(\delta,\mu)$ is strongly continuous by the Banach-Steinhaus Theorem.
To prove (98), consider the inner product for given $\{d_{0,m}\},\{d_{1,m}\}\in\ell^{2}$:
$\left|((\widetilde{\mathbf{H}}(\delta,\mu_{1})-\widetilde{\mathbf{H}}(\delta,\mu_{2}))\{d_{0,m}\},\{d_{1,m}\})_{\ell^{2}}\right|$
$=$ $\sum_{j=1}^{\infty}{\bf i}\left|(\alpha_{j}(\mu_{1})-\alpha_{j}(\mu_{2}))\right|(\phi_{j},\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta)}{\sqrt{1+\lambda_{m}(\delta)}})_{\Gamma_{h}}(\phi_{j},\sum_{m=0}^{\infty}\frac{d_{1,m}\psi_{m}(\cdot,\delta)}{\sqrt{1+\lambda_{m}(\delta)}})_{\Gamma_{h}}$
(106) $\leq$ $C_{3}|\mu_{1}-\mu_{2}|\|\{d_{0,m}\}\|_{\ell^{2}}\|\{d_{1,m}\}\|_{\ell^{2}},$
where we have used the same techniques as in Lemma 4.3 and the assumption $h\ll 1$.
Theorem 5.11.
Under Conditions A.1, A.2, and the assumption $h\ll 1$, the functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are continuous in $\delta$ and Lipschitz continuous in $\mu$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, the following Lipschitz bounds hold:
(107) $|f_{0}(\delta,\mu_{1},h)-f_{0}(\delta,\mu_{2},h)|\leq$ ${C_{f_{0}}}|\mu_{1}-\mu_{2}|,$
(108) $|f_{1}(\delta,\mu_{1},h)-f_{1}(\delta,\mu_{2},h)|\leq$ ${C_{f_{1}}}|\mu_{1}-\mu_{2}|,$
where $C_{f_{0}}<1$ and $C_{f_{1}}<1$ are two constants independent of $\delta$.
Proof 5.12.
By neglecting the diagonal matrix $\mathrm{Diag}(\{\lambda_{m}-\mu\}_{m=0}^{M-1})$, we establish that each entry of the finite-dimensional system (56)–(57) is continuous in $\delta$ and Lipschitz continuous in $\mu$, with a Lipschitz constant of order $\mathcal{O}(h)$. For given indices $0{\leq}m_{1},m_{2}\leq{M-1}$, the entry in the $m_{1}$-th row and $m_{2}$-th column of the matrix $\mathbf{A}$ satisfies
$\Bigg{|}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta_{1}))_{\Gamma_{h}}(\psi_{m_{2}}(\cdot,\delta_{1}),\phi_{j})_{\Gamma_{h}}$
$\qquad\qquad-\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta_{2}))_{\Gamma_{h}}(\psi_{m_{2}}(\cdot,\delta_{2}),\phi_{j})_{\Gamma_{h}}\Bigg{|}$
$\leq$ $n_{1}^{2}\|\frac{1}{n(\cdot,\delta_{1})}\nabla(\psi_{m_{1}}(\cdot,\delta_{1})-\psi_{m_{1}}(\cdot,\delta_{2}))\|_{[L^{2}(R_{h})]^{2}}\|\frac{1}{n(\cdot,\delta_{1})}\nabla(\psi_{m_{2}}(\cdot,\delta_{1}))\|_{[L^{2}(R_{h})]^{2}}$
$+n_{1}^{2}\|\frac{1}{n(\cdot,\delta_{1})}\nabla(\psi_{m_{2}}(\cdot,\delta_{1})-\psi_{m_{2}}(\cdot,\delta_{2}))\|_{[L^{2}(R_{h})]^{2}}\|\frac{1}{n(\cdot,\delta_{2})}\nabla(\psi_{m_{1}}(\cdot,\delta_{2}))\|_{[L^{2}(R_{h})]^{2}}.$
Hence Lemma 3.1 gives its continuity in $\delta$. Moreover, it can be verified that
(109) $\left|\sum_{j=1}^{\infty}{\bf i}(\alpha_{j}(\mu_{1})-\alpha_{j}(\mu_{2}))(\phi_{j},\psi_{m}(\cdot,\delta))_{\Gamma_{h}}(\psi_{m}(\cdot,\delta),\phi_{j})_{\Gamma_{h}}\right|\leq\mathcal{O}(h)|\mu_{1}-\mu_{2}|,$
where the constant hided in $\mathcal{O}$ is irrelevant to $\delta$. Let $\mathbf{z}_{m_{1}}(\delta,\mu)$ denote the vector
(110) $\mathbf{z}_{m_{1}}(\delta,\mu):=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta))_{\Gamma_{h}}\mathbf{w}_{j}(\delta),$
such that $\mathbf{z}_{m_{1}}(\delta,\mu)/\sqrt{1+\lambda_{m_{1}}(\delta)}$ is a restriction of the $m_{1}$-th column of the operator $\widetilde{\mathbf{H}}(\delta,\mu)$ defined as (99). Hence, $\mathbf{z}_{m_{1}}(\delta,\mu)$ is continuous in $\delta$ and Lipschitz continuous in $\mu$. Furthermore, using Lemma 5.1 and the general Sobolev inequalities, the following estimates hold:
(111) $\|\mathbf{z}_{m_{1}}(\delta,\mu)\|_{\ell^{2}}$ $=\mathcal{O}(h),$
(112) $|\mathbf{z}_{m_{1}}(\delta,\mu_{1})-\mathbf{z}_{m_{1}}(\delta,\mu_{2})|$ $=\mathcal{O}(h)|\mu_{1}-\mu_{2}|.$
By similar techniques used in Lemma 5.9, we can show that the vector $\mathbf{w}_{0}(\delta)$ is also continuous in $\delta$. Finally, observe that the $(m_{1},m_{2})$-entry of $\mathbf{V}^{*}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}$ admits the representation $\mathbf{z}_{m_{1}}^{*}\mathbf{T}(\mathbf{I}-\mathbf{T}\mathbf{H}\mathbf{T})^{-1}\mathbf{T}\mathbf{z}_{m_{2}}$, and the $m_{2}$-th column entry of $\mathbf{v}^{*}_{0,M-1,0}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}$ can be written as $\mathbf{w}_{0}^{*}\mathbf{T}(\mathbf{I}-\mathbf{T}\mathbf{H}\mathbf{T})^{-1}\mathbf{T}\mathbf{z}_{m_{2}}$. The desired results follow from Lemmas 5.5, 5.7, and 5.9.
5.3 Existence of Friedrich-Wintgen BICs
In the following theorem, we prove that each equation (72)–(73) defines an implicit function of $\mu(\delta)$. Consequently, BICs correspond to the intersections of the graphs of these functions.
Theorem 5.13.
Under Conditions A.1–A.3 and the assumption $h\ll 1$, the system (3)–(4) admits at least one BIC characterized by $\delta\approx 0$ and $\mu\approx\lambda_{M-1}(0)$.
Proof 5.14.
We restrict $\delta$ and $\mu$ to the intervals $[\delta_{l},\delta_{r}]$ and $[\mu_{l},\mu_{r}]$ respectively, where
(113) $\delta_{l}<0<\delta_{r},\ \mu_{l}=\lambda_{M-3}(0)+2\epsilon,\ \text{and}\ \mu_{r}=\lambda_{M}(0)-2\epsilon.$
Assume $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are continuous in $\delta$ and $\mu$ over these intervals, and (31), (32), (37), (79), (107) and (108) are satisfied. For $h\ll 1$ with $|f_{0}(\cdot,\cdot,h)|<\epsilon$, the following inequalities hold for each fixed $\delta$:
(114) $\lambda_{M-2}(\delta)-\mu_{l}+f_{0}(\delta,\mu_{l},h)>0,$
(115) $\lambda_{M-2}(\delta)-\mu_{r}+f_{0}(\delta,\mu_{r},h)<0.$
By the Intermediate Value Theorem, there exists a unique $\mu_{0}(\delta)\in(\mu_{l},\mu_{r})$ satisfying
(116) $\lambda_{M-2}(\delta)-\mu_{0}(\delta)+f_{0}(\delta,\mu_{0}(\delta),h)=0,$
where the uniqueness follows from the inequality (107). Analyzing the difference between $\mu_{0}(\delta_{1})$ and $\mu_{0}(\delta_{2})$ for $\delta_{1},\delta_{2}\in[\delta_{l},\delta_{r}]$, we obtain
$|\mu_{0}(\delta_{1})-\mu_{0}(\delta_{2})|=$ $|\lambda_{M-2}(\delta_{1})-\lambda_{M-2}(\delta_{2})+f_{0}(\delta_{1},\mu_{0}(\delta_{1}),h)-f_{0}(\delta_{2},\mu_{0}(\delta_{2}),h)|$
$\leq$ $|\lambda_{M-2}(\delta_{1})-\lambda_{M-2}(\delta_{2})|+|f_{0}(\delta_{1},\mu_{0}(\delta_{2}),h)-f_{0}(\delta_{2},\mu_{0}(\delta_{2}),h)|$
$+|f_{0}(\delta_{1},\mu_{0}(\delta_{1}),h)-f_{0}(\delta_{1},\mu_{0}(\delta_{2}),h)|$
$\leq$ $|\lambda_{M-2}(\delta_{1})-\lambda_{M-2}(\delta_{2})|+|f_{0}(\delta_{1},\mu_{0}(\delta_{2}),h)-f_{0}(\delta_{2},\mu_{0}(\delta_{2}),h)|$
(117) $+C_{f_{0}}|\mu_{0}(\delta_{1})-\mu_{0}(\delta_{2})|.$
This implies the continuity of $\mu_{0}(\delta)$ since $C_{f_{0}}<1$. Similarly, there exists a unique continuous function $\mu_{1}(\delta)$ satisfying
(118) $\lambda_{M-1}(\delta)-\mu_{1}(\delta)+f_{1}(\delta,\mu_{1}(\delta),h)=0.$
Without loss of generality, we assume $\lambda_{M-2}(\delta_{r})>\lambda_{M-1}(\delta_{r})$, $\lambda_{M-2}(\delta_{l})<\lambda_{M-1}(\delta_{l})$ and $h\ll 1$ such that
(119) $|f_{0}(\cdot,\cdot,h)|+|f_{1}(\cdot,\cdot,h)|\leq\min\Big{\{}\lambda_{M-2}(\delta_{r})-\lambda_{M-1}(\delta_{r}),\lambda_{M-1}(\delta_{l})-\lambda_{M-2}(\delta_{l})\Big{\}}.$
From this bound, we deduce
(120) $\mu_{0}(\delta_{r})-\mu_{1}(\delta_{r})\geq 0,\ \mu_{0}(\delta_{l})-\mu_{1}(\delta_{l})\leq 0.$
By the Intermediate Value Theorem, $\mu_{0}(\delta)$ and $\mu_{1}(\delta)$ must coincide for at least one $\delta$.
Remark 5.15.
In our analysis, Condition A.1 ensures the continuous dependence of eigenvalues and eigenfunctions on the parameter $\delta$ for the boundary value problem (5)–(6). Specifically, the analytic dependence of the refractive index $n(\cdot,\delta)$ on $\delta$ implies:
  • The continuity of eigenvalues (9),
  • The $H^{1}(\Omega_{\mathrm{in}})$-continuity of eigenfunctions (8).
These results enable the reduction of the infinite-dimensional system (28) to a finite-dimensional system (71) for $\delta$ and $\mu$ over small intervals and $h\ll 1$. By further requiring $n(\mathbf{x},\delta)=1$ for $\mathbf{x}\in\Omega_{B}$, we can derive the uniform boundedness of the eigenfunctions and their derivatives on $\Gamma_{h}$ and its neighboring regions as $\delta$ varies. Combining Condition A.1 with Condition A.2, the two governing equations (72)–(73) are well-defined. Furthermore, these conditions guarantee the following critical properties for the functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ defined in (74)–(75) for $h\ll 1$:
  • uniform convergence: $f_{0},f_{1}\to 0$ as $h\to 0$ uniformly;
  • continuity in $\delta$: $f_{0}$ and $f_{1}$ are continuous with respect to $\delta$;
  • Lipschitz continuity in $\mu$: both $f_{0}$ and $f_{1}$ are Lipschitz continuous in $\mu$ with a Lipschitz constant less than 1.
Consequently, each governing equations (72)–(73) defines an implicit function $\mu(\delta)$. BICs emerge as intersections of these implicit functions. Under Condition A.3, the existence of such an intersection is rigorously guaranteed.
6 Numerical experiments
In [16], the emergence of FW-BICs in an acoustic system was numerically demonstrated for homogeneous rectangular cavities with waveguide openings. If the cavity supports a multiple eigenvalue, a FW-BIC can be achieved by perturbing the cavity length slightly. Here, we consider an inhomogeneous, geometrically asymmetric region, governed by the PDE system (3)–(4). Through numerical experiments, we show that a FW-BIC can be achieved by introducing a slight perturbation to the refractive index $n(x)$, assuming the closed cavity has a multiple eigenvalue when there is no perturbation.
As illustrated in Fig. 2, the full computational region $\Omega$ is partitioned into three subregions:
  • $\Omega_{\mathrm{in}}$: the rectangular cavity,
  • $\Omega_{\mathrm{out}}$: a semi-infinite waveguide of width $h=2\pi/9$,
  • $\Gamma_{h}$: the shared boundary between $\Omega_{\mathrm{in}}$ and $\Omega_{\mathrm{out}}$.
A rectangular coordinate system is centered at the midpoint $\mathbf{o}$ of $\Gamma_{h}$, with the cavity extending between the coordinates $\mathbf{o}_{1}=(0,-4\pi/3)$ and $\mathbf{o}_{2}=(-\pi,2\pi/3)$. Six identical circular elements (radius 0.48) are symmetrically embedded within the cavity. Their centers are located at:
(121) $(-\pi/2\pm 0.8,\pi/3),\ (-\pi/2\pm 0.8,-\pi/3)\ \text{and}\ (-\pi/2\pm 0.8,-\pi).$
The refractive index $n(\mathbf{x})$ is modeled as a piecewise constant function: it assumes the same constant value within each element and is set to $1$ in the surrounding ambient medium.
$\mathbf{o}$$x_{1}$$x_{2}$$\Omega_{\mathrm{out}}$$\Omega_{\mathrm{in}}$$n(\mathbf{x})$$\mathbf{o}_{1}$$\mathbf{o}_{2}$$\Omega_{\mathrm{out}}$$\Gamma_{h}$$h$
Figure 2: Schematic of a rectangular cavity with a waveguide opening. The computational region is defined as $\Omega:=\Omega_{\mathrm{in}}\cup\Gamma_{h}\cup\Omega_{\mathrm{out}}$ comprises a rectangular cavity $\Omega_{\mathrm{in}}$ spanning between coordinates $\mathbf{o}_{1}=(0,-4\pi/3)$ and $\mathbf{o}_{2}=(-\pi,2\pi/3)$, a semi-infinite waveguide $\Omega_{\mathrm{out}}$ of width $h=2\pi/9$, and their shared interface $\Gamma_{h}$. A rectangular coordinate system is centered at $\mathbf{o}$, the midpoint of $\Gamma_{h}$. Six identical circular elements (radius 0.48) with refractive index $n(\cdot)$ are symmetrically embedded in $\Omega_{\mathrm{in}}$, centered at positions $(-\pi/2\pm 0.8,\pi/3)$, $(-\pi/2\pm 0.8,-\pi/3)$ and $(-\pi/2\pm 0.8,-\pi)$.
For the eigenvalue problem (5)–(6) in the region $\Omega_{\mathrm{in}}$, Figure 3 illustrates the evolution of the first eight eigenvalues as the refractive index $n$ of the elements increases from $1$ to $2$. Notably, the sixth and seventh eigenvalues, labeled $\lambda_{6}$ and $\lambda_{7}$, cross at $n\approx 1.461$, resulting in a multiple eigenvalue $\lambda_{6}=\lambda_{7}\approx 1.695$.
When a thin waveguide is introduced on the boundary of $\Omega_{\mathrm{in}}$, the resulting resonances approximating $\lambda_{6}$ and $\lambda_{7}$ are denoted by $\widetilde{\lambda}_{6}$ and $\widetilde{\lambda}_{7}$. Their real and imaginary parts are plotted in Fig. 4(a) and 4(b), respectively. A BIC emerges at $n\approx 1.442$ and $\widetilde{\lambda}_{6}\approx 1.718$, confirming our theoretical predictions.
Figure 3: First eight eigenvalues of the problem (5)–(6) in $\Omega_{\mathrm{in}}$ as the refractive index of the elements varies from $1$ to $2$.
(a) Real parts of $\widetilde{\lambda}_{6}$ and $\widetilde{\lambda}_{7}$.
(b) Imaginary parts of $\widetilde{\lambda}_{6}$ and $\widetilde{\lambda}_{7}$.
Figure 4: Resonances of the problem (3)–(4) in $\Omega$ approximating the eigenvalues $\lambda_{6}$ and $\lambda_{7}$ as the refractive index of the elements varies.
7 Discussion and conclusion
In this section, we briefly discuss how to extend our approach to analyze the existence of FW-BICs under generalized perturbations, including parameter-dependent boundary perturbations. Additionally, the emergence of symmetry-protected BICs in the cavity-waveguide structures are also investigated.
Previously, we assumed the refractive index $n(\mathbf{x},\delta)\equiv 1$ within $\Omega_{B}$. If $n(\mathbf{x},\delta)$ also exhibits dependence on $\delta$ in $\Omega_{B}$, our approach remains applicable with minor modifications. This requires the following conditions:
  • B.1.
    $n(\mathbf{x},\delta)|_{{\overline{\Omega_{B}}}}\in{C^{1,1}(\overline{\Omega_{B}}\times[-\delta_{0},\delta_{0}])}$,
  • B.2.
    $n(\mathbf{x},\delta)|_{\Gamma_{h}}=C_{\delta}$, where $C_{\delta}$ is a constant for each $\delta$.
Under such perturbations, the second Green identity (25) becomes:
(122) $(\mathcal{P}u,\psi_{m})_{\Omega_{\mathrm{in}}}-(u,\mathcal{P}\psi_{m})_{\Omega_{\mathrm{in}}}=(u,\frac{1}{n^{2}}\partial_{\bm{\nu}}\psi_{m})_{\partial\Omega_{\mathrm{in}}}-(\frac{1}{n^{2}}\partial_{\bm{\nu}}u,\psi_{m})_{\partial\Omega_{\mathrm{in}}}.$
Condition B.2 ensures the simplification:
(123) $(\frac{1}{n^{2}}\partial_{\bm{\nu}}u,\psi_{m})_{\partial\Omega_{\mathrm{in}}}=\frac{1}{C_{\delta}^{2}}(\partial_{\bm{\nu}}u,\psi_{m})_{\Gamma_{h}}$
by leveraging the boundary condition (4). Our reduction techniques then apply directly to the infinite-dimensional linear system:
(124) $\left(\begin{bmatrix}\lambda_{0}-\mu&&\\ &\lambda_{1}-\mu&\\ &&\ddots\end{bmatrix}-\sum_{j=0}^{\infty}\frac{{\bf i}\alpha_{j}}{C_{\delta}^{2}}\mathbf{v}_{j}\mathbf{v}^{*}_{j}\right)\begin{bmatrix}d_{0}\\ d_{1}\\ \vdots\end{bmatrix}=\mathbf{0}.$
By imposing Condition B.1, we retain the eigenfunction regularity proven in Lemma 5.1. Consequently, the existence of FW-BICs analogously follows the analysis presented in Section 5.
Our approach can also be generalized to linear boundary perturbations governed by one-to-one transformations:
[…]
The scaling transformation considered in [16] is equivalent to defining $\bm{\mathcal{R}}$ as:
[…]
Letting $\Gamma$ denote the line segment on $\partial\Omega_{\mathrm{in}}$ where the waveguides extend from, we require
  • R.1.
[…]
  • R.2.
    The gradient of $\bm{\mathcal{R}}$ satisfies
[…]
    where $C_{\bm{\mathcal{R}}}$ is a constant.
Under Condition R.2, $\bm{\mathcal{T}}_{\delta}(\Gamma)$ is just a translation of $\Gamma$. Hence, our perturbed cavity-waveguide region $\Omega_{\delta}$ is well-defined by assuming the waveguide translates along with $\Gamma$. For simplicity, we assume $n(\mathbf{x})\equiv 1$ in $\Omega$. Letting $\mathbf{S}_{\delta}:=\nabla(\bm{\mathcal{T}}_{\delta}^{-1})$ and $J_{\delta}:=\mathrm{det}(\nabla\bm{\mathcal{T}}_{\delta})$, solving the PDE problem (3)–(4) in $\Omega_{\delta}$ is equivalent to solving the following equations in $\Omega$:
(128) $-\nabla\cdot\left(\mathbf{S}_{\delta}^{T}J_{\delta}\mathbf{S}_{\delta}\nabla{u}\right)+\mu{J_{\delta}}u=$ $0\ \text{in}\ \Omega_{\mathrm{in}},$
(129) $\bm{\nu}\cdot\mathbf{S}_{\delta}^{T}J_{\delta}\mathbf{S}_{\delta}{\nabla}u=$ $0\ \text{on}\ \partial\Omega_{\mathrm{in}}/\Gamma_{h},$
(130) $-\Delta{u}+\mu{u}=$ $0\ \text{in}\ \Omega_{\mathrm{out}},$
(131) $\partial_{\bm{\nu}}u=$ $0\ \text{on}\ \partial\Omega_{\mathrm{out}}/\Gamma_{h},$
(132) $u|_{\Gamma_{h,-}}=$ $u|_{\Gamma_{h,+}},$
(133) $\bm{\nu}\cdot\mathbf{S}_{\delta}^{T}J_{\delta}\mathbf{S}_{\delta}{\nabla}u|_{\Gamma_{h,-}}=$ $\partial_{x_{1}}u|_{\Gamma_{h,+}},$
where $\Gamma_{h,-}$ and $\Gamma_{h,+}$ denote the sides of $\Gamma_{h}$ approached from $\Omega_{\mathrm{in}}$ and $\Omega_{\mathrm{out}}$ respectively. In the $\Omega_{\mathrm{in}}$ region, the eigenfunction expansion remains valid by considering the following generalized eigenvalue problem:
(134) $-\nabla\cdot\left(\mathbf{S}_{\delta}^{T}J_{\delta}\mathbf{S}_{\delta}\nabla{\psi}\right)+\lambda{J_{\delta}}\psi=$ $0\ \text{in}\ \Omega_{\mathrm{in}},$
(135) $\bm{\nu}\cdot\mathbf{S}_{\delta}^{T}J_{\delta}\mathbf{S}_{\delta}{\nabla}\psi=$ $0\ \text{on}\ \partial\Omega_{\mathrm{in}},$
where the eigenfunctions form a complete orthonormal basis in $L^{2}(\Omega_{\mathrm{in}})$ with the weight $J_{\delta}$. The continuity of the eigenvalues and eigenfunctions as $\delta$ varies can also be established under the transformation (125) and Condition R.1. (cf. Section VII.6.5 in [10]). By applying a similar mode-matching method and invoking Condition R.2, we derive the following infinite-dimensional linear system:
[…]
By requiring that Condition A.2 and A.3 are satisfied, the existence of FW-BICs can be established via an argument analogous to the analysis presented in Section 5.
Symmetry-protected BICs are a class of BICs frequently observed in symmetric structures. Their existence arises from a symmetry mismatch between the bound states and the outgoing radiating waves [1, 2, 25]. The mode-matching method can be used to rigorously establish the existence of symmetry-protected BICs in cavity-waveguide structures, if the waveguide widths are sufficiently small. Assuming the region $\Omega$ is symmetric with respect to $x_{2}=0$, the system (28)–(29) decouples into two subsystems governed by eigenfunctions with the following symmetries:
  • even: $\psi^{(e)}(x_{1},x_{2})=\psi^{(e)}(x_{1},-x_{2})$,
  • odd: $\psi^{(o)}(x_{1},x_{2})=-\psi^{(o)}(x_{1},-x_{2})$.
For the odd subsystem, (41) is automatically satisfied. Following the reduction techniques outlined in Section 4, we obtain the governing equation:
(137) $\lambda-\mu+a(\mu,h)=0,$
where $\lambda$ is a simple eigenvalue associated with an odd eigenfunction and $a(\mu,h)$ represents a function depending on $\mu$ and $h$. By demonstrating the uniform convergence of $a(\cdot,h)$ to 00 as $h\to 0$ and the continuity of $a(\mu,h)$ in $\mu$, the existence of a BIC is guaranteed for $h\ll 1$ by the Intermediate Value Theorem.